# Vulnerability Webs:

### v2

@@ 2.2 Dependency Graphs
Most open source software is organized by package managers like PyPi for the Python programming language that standardize and automate software distribution. To make the installation of complex software projects easier for users, package managers keep track of a package’s dependencies and allow users to install these alongside the package they are installing. The existence of package managers and the level of automation they provide is necessary because modern software is frequently updated and different versions of the same package are not always compatible. Packages are logically grouped into repositories, controlled and managed by a maintainer. [^15] Maintainers can also be organizations. Some repositories are also controlled by a group of maintainers, but for our purposes, this is immaterial.
@@
Table 2: Distribution of in- and out-degree for the largest weakly connected component of the repo-based dependency graph of the PyPi ecosystem with $N=16,102$ nodes.
@@
In Table LABEL:Tab::TopRepos we show the $10$ most depended-on packages (i.e. with the highest in-degree) The repository with highest in-degree is six, which facilitates compatibility between python version 2 and 3 (where some major updates have led to serious incompatibility issues). The second-highest in-degree repository is pytest-cov, which is used to check a code’s test coverage, which is a crucial step in modern software development. Likewise, the third-most commonly linked software package is mock, which is also used for software testing.
@@
The largest weakly connected component of the repo-based dependency graph is shown in Figure LABEL:Fig::RustDependencyGraph, where we apply the community detection algorithm of Blondel et al. (2008) to color-code packages in the same community. The three largest communities contain about $10-11\%$ of all repositories each.
@@ 2.3 Observables: Covariates
In addition to the software dependency data, we obtain additional data from the from libraries.io website for each package (e.g. https://libraries.io/pypi/pandas), which we show for the ten most-depended on repositories in Table LABEL:Tab::TopRepos. We use three types of covariates, summarized in a vector of observable attributes $\bm{x}_{i}$. First, we use a repository’s Size of all code in the repository, measured in Kilobytes. Second, Popularity is measured as the number of stars–an expression of how many people like a repository or find it useful–a repository received on the website hosting it, e.g., GitHub. And, third, the time in days between the first release of a repository and the reference date of 13 January 2020 is denoted Maturity.
@@
Table 4: Distribution of covariates
To facilitate the estimation of our model, we create a categorical variable for each of our covariates using quartiles of the distribution. The reason for discretizing variables is computational, as our algorithm is based on stochastic block models with discrete types (Bickel et al., 2013; Vu et al., 2013). If the covariates are discrete, the estimation engine of stochastic block models can be adapted to estimate the unobserved heterogeneity (Vu et al., 2013; Babkin et al., 2020; Dahbura et al., 2021). Otherwise, when the covariates are continuous, the computational costs are prohibitive for such large networks. [^21] We provide more details in the appendix; see also Babkin et al. (2020) and Dahbura et al. (2021).
@@ 3 Model
We model software development as a dynamic process, assuming that the developers update their packages to add or improve their functionality. Maintainers can update their packages by either adding functionalities to their own packages or by creating dependencies to other packages that have the desired functionalities.
@@ 3.1 Setup
Each package $i$ and its developer team is characterized by a vector of $P$ observable attributes $\bm{x}_{i}\in\mathbb{R}^{P}$, which are stored in a matrix $\bm{x}\in\mathbb{R}^{N\times P}$. We assume that each of the $N$ packages belongs to one of $K\geq 2$ types, unknown to researchers but known to other developers. The type of a software package may be its purpose, which may be known to developers but may not be known to researchers, or may correspond to unobservable qualities of the code and its developer team. To indicate the type $k$ of package $i$, we define $z_{ik}\coloneqq 1$ and $z_{il}\coloneqq 0$ for all $l\neq k$ if package $i$ is of type $k$, and write $\bm{z}_{i}\coloneqq(z_{i1},\ldots,z_{iK})$. We assume that the unobservables are independent and identically distributed according to a multinomial distribution:
@@
where $\bm{\eta}\coloneqq(\eta_{1},\ldots,\eta_{K})$ is a vector of $K$ probabilities summing to $1$. We represent the software dependency network $\bm{g}$ by the adjacency matrix, with elements $g_{ij}=1$ if package $i$ depends on package $j$ and $g_{ij}=0$ otherwise. In general, software dependency networks are directed: If package $i$ depends on package $j$, then package $j$ need not depend on package $i$.
The utility that developers of package $i$ receive from the software dependency network $\bm{g}$, with observables $\bm{x}$, unobservables $\bm{z}$, and parameters $\bm{\theta}\coloneqq(\bm{\alpha},\,\bm{\beta},\,\gamma)$, is of the form
@@
The function $u_{ij}(\bm{\alpha},\,\bm{\beta})\coloneqq u(\bm{x}_{i},\,\bm{x}_{j},\bm{z}_{i},\,\bm{z}_{j},\bm{\alpha},\,\bm{\beta})$ is the direct utility of linking package $i$ to package $j$. It is a function of observables ($\bm{x}_{i},\,\bm{x}_{j}$), unobservables ($\bm{z}_{i},\,\bm{z}_{j}$), and parameters ($\bm{\alpha},\,\bm{\beta}$). This represents the benefit of a dependency of package $i$ on package $j$, net of costs: e.g., the developers will have to audit the code of package $j$ and determine its quality, and have to maintain the dependency on package $j$. The cost may includes modification and adaptation of the code, because the developers have to be able to use the functions and methods available in package $j$. Additional costs may arise from relying on other packages, because other packages may contain undetected bugs. Finally, relying on existing code is convenient but may decrease the incentives of developers to maintain and improve coding skills, which in turn can undermine the future productivity of developers and can hence be viewed as a cost.
Motivated by the application to software dependency networks, we parameterize $u_{ij}$ as follows:
$u(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha},\,\bm{\beta})\;\coloneqq\;\begin{cases}u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})\ &\text{ if both $i$ and $j$ are of type $k$}\vspace{.25cm}\\ u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})\ &\text{otherwise,}\end{cases}$ (3)
where the utility for dependencies within types $u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})$ is
$\begin{array}[]{lllll}u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})&\coloneqq&\alpha_{w1}+\alpha_{w1}^{\ell}\,\log(N_{k})+\displaystyle\sum_{p=1}^{P}\beta_{wp}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j})\\ &+&\displaystyle\sum_{p=1}^{P}\beta_{wp}^{\ell}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j})\,\log(N_{k}),\end{array}$ (4)
and the utility for dependencies between types $u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})$ is
$\begin{array}[]{lllll}u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})&\coloneqq&\alpha_{b}+\displaystyle\sum_{p=1}^{P}\beta_{bp}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j}).\end{array}$ (5)
@@
The terms in the within- and between-type utilities can be interpreted as follows: The intercept $\alpha_{w1}\in\mathbb{R}$ represents the fixed cost of forming a dependency; the term $\alpha_{w2}\,\log(N_{k})$ with weight $\alpha_{w2}\in\mathbb{R}$ is the part of the cost that depends on the size $N_{k}$ of type $k$; [^22] In principle, the specification can be extended to incorporate aggregated terms at the type-level. In this sense, $\log(N_{k})$ is one among other examples of terms that can be used to make the model more flexible. For instance, one may include the average of observed characteristics of type $k$ as one of the covariates. We do not pursue this specification here, given the very sparse nature of the software dependency network. and $h_{ijp}(\bm{x}_{i},\bm{x}_{j})\coloneqq\bm{1}\{{x}_{ip}={x}_{jp}\}$ is an indicator function, which is equal to $1$ if the $p$th covariate of $i$ and $j$ matches and is 00 otherwise. Size-dependent terms of the form $\log(N)$ were suggested by Krivitsky et al. (2011), motivated by invariance considerations (encouraging the expected number of links to be invariant to $N$), and adapted to observed types by Krivitsky et al. (2023) and to unobserved types by Babkin et al. (2020). The micro-behavioral foundations of size-dependent terms of the form $\log(N)$ were studied by Butts (2019). The vectors $\bm{\alpha}\coloneqq(\alpha_{w1},\,\alpha_{w2},\,\alpha_{b})\in\mathbb{R}^{3}$ and $\bm{\beta}\coloneqq(\beta_{w1},\,\ldots,\,\beta_{wP},\,\beta_{b1},\,\ldots,\,\beta_{bP},\beta_{w1}^{\ell},\,\ldots,\,\beta_{wP}^{\ell})\in\mathbb{R}^{3P}$ are parameters to be estimated.
@@
The second and third term of utility function (2) are externalities generated by linking to package $j$, because package $j$ may be linked to other packages $r$ as well. This means that the developers of package $i$ need to check the quality and features of the code in those packages $r$ as well, to ensure that those packages are free of bugs or vulnerabilities, and compatible with package $i$. On the other hand, any update of package $r$ may compromise compatibility with packages $i$ and $j$, increasing the cost of maintaining package $i$. In short, $w_{ijr}$ measures the net benefits of this externality when developers of package $i$ form a dependency to package $j$. In addition, if the developers of package $i$ create a dependency on package $j$, it may compromise compatibility with other packages $r$ that depend on package $i$. This is accounted for by the third term $v_{ijr}$ in utility function (2). We assume that the second and third term in the utility are of the form
@@
The choice of the utility functions $U_{i}$ guarantees that all externalities are local, in the sense that all externalities are limited to packages of the same type. Local externalities make sense in large networks, because developers are unlikely to know the functionalities of thousands of other packages and, more likely than not, will create dependencies to other packages of the same type, which developers are most familiar with. In addition, local externalities have probabilistic, statistical, and computational advantages. First, local externalities ensure desirable properties of models for large networks (Schweinberger and Handcock, 2015; Schweinberger et al., 2020). Second, local externalities facilitate theoretical guarantees for statistical procedures (Schweinberger and Stewart, 2020; Schweinberger, 2020; Stewart and Schweinberger, 2023; Stewart, 2024). Third, local externalities enable local computing on subnetworks of large networks, which facilitates parallel computing on a massive scale (Babkin et al., 2020).
@@ 3.2 Equilibrium
Software development is a dynamic process. We assume that time is discrete and that at each time point $t=1,2,\ldots$ a single package is updated. In other words, maintainers do not update software packages continuously, but certain events – unobservable to researchers – trigger the need for an update. For example, if package $i$ depends on package $j$, then an update of package $j$ may require an update of package $i$, or a request by users of package $i$ may trigger an update of package $i$. The update can be produced in-house or by forming a dependency to some other package $j$ containing the same functionality. Since researchers may not be able to observe the timing of such updates, we model updates as a discrete-time Markov process (Norris, 1997).
@@
The first ingredient if the discrete-time Markov process is the update process that determines which package $i$ requires an update and which dependency $j$ is considered to implement the update. We assume that the probability that package $i$ requires an update and the developers of package $i$ consider creating, maintaining, or removing a dependence to package $j$ is $\rho(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j})$, where $\bm{g}^{t-1}$ is the network at time $t-1$. So the update process may depend on the existing network $\bm{g}^{t-1}$ as well as observables ($\bm{x}_{i}$, $\bm{x}_{j}$) and unobservables ($\bm{z}_{i}$, $\bm{z}_{j}$). A simple example is an update process that selects an ordered pair of packages $i$ and $j$ at random. In other words, a package $i$ is selected with probability $1/n$ and requires an update. Conditional on the event that package $i$ requires an update, the developers of package $i$ consider updating the possible dependency of package $i$ on package $j$ with probability $1/(n-1)$. As a result, the probability that package $i$ requires an update and the developers of package $i$ consider updating a (possible) dependency of package $i$ on package $j$ is
$\rho\left(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)\;\;=\;\;\dfrac{1}{n(n-1)}.$
A second example is an update process that allows the conditional probability that an update of the (possible) dependency of package $i$ on package $j$ is proposed to depend on the distance $d(x_{i},x_{j})$ between packages $i$ and $j$:
$\rho\left(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)\;\;\propto\;\;\frac{1}{n}\exp\left(-d(\bm{x}_{i},\bm{x}_{j})\right),$
where $d(\bm{x}_{i},\bm{x}_{j})$ is a function of observables ($\bm{x}_{i}$, $\bm{x}_{j}$) satisfying reflexity, symmetry, and the triangle inequality. In other words, $d(\bm{x}_{i},\bm{x}_{j})$ is a distance function that quantifies the distance between packages $i$ and $j$ based on observable characteristics $\bm{x}_{i}$ and $\bm{x}_{j}$ of packages $i$ and $j$.
@@
$\rho\Big{(}\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\Big{)}\;\;=\;\;\rho\Big{(}\bm{g}_{-(i,j)}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\Big{)}.$
Second, the update probabilities $\rho(\bm{g}_{-(i,j)}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j})$ are strictly positive for all ordered pairs of packages $(i,j)$:
[…]
This implies that each possible dependency can be proposed with a positive probability $\rho$, no matter how small. These assumptions help ensure that in the long-run equilibrium these probabilities do not affect the distribution of networks (Mele, 2017).
While the update process is stochastic, the decision to create, maintain, or remove a possible dependency of package $i$ on package $j$ is a strategic decision. Conditional on the opportunity to update the dependency of package $i$ on package $j$, the developers of package $i$ make a decision about whether to create, maintain, or remove a dependency on package $j$ based on utility function $U_{i}(\bm{g};\,\bm{x},\bm{z},\bm{\theta})$. Before the developers of package $i$ make a decision about the (possible) dependency on package $j$, the developers receive random shocks $\varepsilon_{ij0}\in\mathbb{R}$ and $\varepsilon_{ij1}\in\mathbb{R}$ to the perceived utility of a dependency of package $i$ on package $j$ ($g_{ij}^{t}=1$) and the perceived utility of writing code in-house ($g_{ij}^{t}=0$). After receiving random shocks $\varepsilon_{ij0}\in\mathbb{R}$ and $\varepsilon_{ij1}\in\mathbb{R}$, the developers of package $i$ create a dependency on package $j$ if the perceived utility of the dependency of package $i$ on package $j$ ($g_{ij}^{t}=1$) exceeds the utility of writing code in-house ($g_{ij}^{t}=0$):
@@
where $\bm{g}_{-(i,j)}^{t-1}$ refers to network $\bm{g}^{t-1}$ excluding the possible dependency $g_{ij}$ of package $i$ on package $j$. The random shocks to payoffs $\varepsilon_{ij0}^{t}\in\mathbb{R}$ and $\varepsilon_{ij1}^{t}\in\mathbb{R}$ capture the fact that unexpected events can occur when developing code in-house ($\varepsilon_{ij0}$) and when creating a dependency ($\varepsilon_{ij1}$), e.g., changes in the composition of the developer team or additional information about package $j$. We assume that the random shocks to payoffs $\varepsilon_{ij0}$ and $\varepsilon_{ij1}$ do not depend on time, are independent across pairs of packages $(i,j)$, and follow a logistic distribution, which is a standard assumption in the literature on discrete choice models (Mele, 2017; Graham and dePaula, 2020; Boucher and Mourifie, 2017). As a consequence, when the opportunity to update the dependency of package $i$ on package $j$ arises, the conditional choice probability of a dependency of package $i$ on package $j$ is
$\mathbb{P}\left(g_{ij}^{t}=1\ |\ \bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta}\right)=\frac{\exp\left(U_{i}(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})-U_{i}(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})\right)}{1+\exp\left(U_{i}(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})-U_{i}(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})\right)}.$ (8)

### v3

@@ 2.2 Dependency Graphs
Most open source software is organized by package managers like PyPi for the Python programming language that standardize and automate software distribution. To make the installation of complex software projects easier for users, package managers keep track of a package’s dependencies and allow users to install these alongside the package they are installing. The existence of package managers and the level of automation they provide is necessary because modern software is frequently updated and different versions of the same package are not always compatible. Packages are logically grouped into repositories, controlled and managed by a maintainer. [^16] Each package belongs to one repository, but a repository can contain multiple packages. Maintainers can also be organizations. Some repositories are also controlled by a group of maintainers, but for our purposes, this is immaterial.
@@
Table 2: In- and out-degree distribution.
@@
In Table LABEL:Tab::TopRepos we show the $10$ most depended-on repositories (i.e. with the highest in-degree). The repository with highest in-degree is six, which facilitates compatibility between python version 2 and 3 (where some major updates have led to serious incompatibility issues). The second-highest in-degree repository is pytest-cov, which is used to check a code’s test coverage, which is a crucial step in modern software development. Likewise, the third-most commonly linked software repository is mock, which is also used for software testing.
@@
The largest weakly connected component of the repo-based dependency graph is shown in Figure LABEL:Fig::RustDependencyGraph, where we apply the community detection algorithm of Blondel et al. (2008) to color-code repositories in the same community. The three largest communities contain about $10-11\%$ of all repositories each.
@@ 2.3 Observables: Covariates
In addition to the software dependency data, we obtain additional data from the libraries.io website for each package and repository (e.g. https://libraries.io/pypi/pandas), which we show for the ten most depended-on repositories in Table LABEL:Tab::TopRepos. We use three types of covariates, summarized in a vector of observable attributes $\bm{x}_{i}$. First, we use a repository’s Size of all code in the repository, measured in Kilobytes. Second, Popularity is measured as the number of stars–an expression of how many people like a repository or find it useful–a repository received on the website hosting it, e.g., GitHub. And, third, the time in days between the first release of a repository and the reference date of 13 January 2020 is denoted Maturity.
@@
Table 4: Distribution of covariates.
To facilitate the estimation of our model, we create a categorical variable for each of our covariates using quartiles of the distribution. The reason for discretizing variables is computational, as our algorithm is based on stochastic block models with discrete types (Bickel et al., 2013; Vu et al., 2013). If the covariates are discrete, the estimation engine of stochastic block models can be adapted to estimate the unobserved heterogeneity (Vu et al., 2013; Babkin et al., 2020; Dahbura et al., 2021). Otherwise, when the covariates are continuous, the computational costs are prohibitive for such large networks.
@@ 3 Model
We model software development as a dynamic process, assuming that maintainers update their repositories to add new functionality or improve existing functionality. Maintainers can update their repositories by adding functionality to their own code or creating dependencies to other repositories that provide the desired functionality.
@@ 3.1 Setup
Each repository $i$ and its maintainer is characterized by a vector of $P$ observable attributes $\bm{x}_{i}\in\mathbb{R}^{P}$, which are stored in a matrix $\bm{x}\in\mathbb{R}^{N\times P}$. We assume that each of the $N$ repositories belongs to one of $K\geq 2$ types, unknown to researchers but known to other developers. The type of a repository may be its purpose, which may be known to developers but may not be known to researchers, or may correspond to unobservable qualities of the code and its developer team. To indicate the type $k$ of repository $i$, we define $z_{ik}\coloneqq 1$ and $z_{il}\coloneqq 0$ for all $l\neq k$ if repository $i$ is of type $k$, and write $\bm{z}_{i}\coloneqq(z_{i1},\ldots,z_{iK})$. We assume that the unobservables are independent and identically distributed according to a multinomial distribution:
@@
where $\bm{\eta}\coloneqq(\eta_{1},\ldots,\eta_{K})$ is a vector of $K$ probabilities summing to $1$. We represent the software dependency network $\bm{g}$ by the adjacency matrix, with elements $g_{ij}=1$ if repository $i$ depends on repository $j$ and $g_{ij}=0$ otherwise. In general, software dependency networks are directed: If repository $i$ depends on repository $j$, then repository $j$ need not depend on repository $i$.
The utility that developers of repository $i$ receive from the software dependency network $\bm{g}$, with observables $\bm{x}$, unobservables $\bm{z}$, and parameters $\bm{\theta}\coloneqq(\bm{\alpha},\,\bm{\beta},\,\gamma)$, is of the form
@@
The function
$u_{ij}(\bm{\alpha},\,\bm{\beta})\;\coloneqq\;u(\bm{x}_{i},\,\bm{x}_{j},\bm{z}_{i},\,\bm{z}_{j},\bm{\alpha},\,\bm{\beta})$ (3)
@@
is the direct utility of linking repository $i$ to repository $j$. It is a function of observables ($\bm{x}_{i},\,\bm{x}_{j}$), unobservables ($\bm{z}_{i},\,\bm{z}_{j}$), and parameters ($\bm{\alpha},\,\bm{\beta}$). This represents the benefit of a dependency of repository $i$ on repository $j$, net of costs: e.g., the developers will have to audit the code of repository $j$ and determine its quality, and have to maintain the dependency on repository $j$. The cost may include modification and adaptation of the code, because developers must be able to use the functions and methods available in the repository $j$. Additional costs may arise from relying on other repositories, because other repositories may contain undetected vulnerabilities. Finally, relying on existing code is convenient but may decrease the incentives of developers to maintain and improve coding skills, which can undermine their future productivity and can hence be viewed as a cost.
Motivated by the application to software dependency networks, we parameterize $u$ as follows:
$u(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha},\,\bm{\beta})\;\coloneqq\;\alpha(\bm{z}_{i},\bm{z}_{j})+\displaystyle\sum_{p=1}^{P}\beta_{p}(\bm{z}_{i},\bm{z}_{j})\,h_{ijp}(\bm{x}_{i},\bm{x}_{j}),$ (4)
where
$\begin{array}[]{lllll}\alpha(\bm{z}_{i},\bm{z}_{j})=\begin{cases}\alpha_{w1}+\alpha_{w1}^{\ell}\,\log(N_{k}),&\text{ if }\bm{z}_{i}=k\text{ and }\bm{z}_{j}=k\\ \alpha_{b},&\text{ otherwise }\\ \end{cases}\end{array}$ (5)
and
$\begin{array}[]{lllll}\beta_{p}(\bm{z}_{i},\bm{z}_{j})=\begin{cases}\beta_{wp}+\beta_{wp}^{\ell}\,\log(N_{k}),&\text{ if }\bm{z}_{i}=k\text{ and }\bm{z}_{j}=k\\ \beta_{bp},&\text{ otherwise }\end{cases}\end{array}$ (6)
define within- and between-type utilities.
@@
The terms in the within- and between-type utilities can be interpreted as follows: The intercept $\alpha_{w1}\in\mathbb{R}$ represents the fixed cost of forming a dependency; the term $\alpha_{w1}^{\ell}\,\log(N_{k})$ with weight $\alpha_{w1}^{\ell}\in\mathbb{R}$ is the part of the cost that depends on the size $N_{k}$ of type $k$; [^22] In principle, the specification can be extended to incorporate aggregated terms at the type-level. In this sense, $\log(N_{k})$ is one among other examples of terms that can be used to make the model more flexible. For instance, one may include the average of observed characteristics of type $k$ as one of the covariates. We do not pursue this specification here, given the very sparse nature of the software dependency network. and $h_{ijp}(\bm{x}_{i},\bm{x}_{j})\coloneqq\bm{1}\{{x}_{ip}={x}_{jp}\}$ is an indicator function, which is equal to $1$ if the $p$th covariate of $i$ and $j$ matches and is 00 otherwise. Size-dependent terms of the form $\log(N)$ were suggested by Krivitsky et al. (2011), motivated by invariance considerations (encouraging the expected number of links to be invariant to $N$), and adapted to observed types by Krivitsky et al. (2023) and to unobserved types by Babkin et al. (2020). The micro-behavioral foundations of size-dependent terms of the form $\log(N)$ were studied by Butts (2019). The vectors $\bm{\alpha}\coloneqq(\alpha_{w1},\,\alpha_{w1}^{\ell},\,\alpha_{b})\in\mathbb{R}^{3}$ and $\bm{\beta}\coloneqq(\beta_{w1},\,\ldots,\,\beta_{wP},\,\beta_{b1},\,\ldots,\,\beta_{bP},\beta_{w1}^{\ell},\,\ldots,\,\beta_{wP}^{\ell})\in\mathbb{R}^{3P}$ are parameters to be estimated.
@@
The second and third term of utility function (2) are externalities generated by linking to repository $j$, because repository $j$ may be linked to other repositories $r$ as well. This means that the developers of repository $i$ need to check the quality and features of the code in those repositories $r$ as well, to ensure that those repositories are free of vulnerabilities, and compatible with repository $i$. On the other hand, any update of repository $r$ may compromise compatibility with repositories $i$ and $j$, increasing the cost of maintaining repository $i$. In short, $w_{ijr}$ measures the net benefits of this externality when developers of repository $i$ form a dependency to repository $j$. In addition, if the developers of repository $i$ create a dependency on repository $j$, it may compromise compatibility with other repositories $r$ that depend on repository $i$. This is accounted for by the third term $v_{ijr}$ in utility function (2). We assume that the second and third term in the utility are of the form
@@
The choice of the utility functions $U_{i}$ guarantees that all externalities are local, in the sense that all externalities are limited to repositories of the same type. Local externalities make sense in large networks, because developers are unlikely to know the functionalities of thousands of other repositories and, more likely than not, will create dependencies to other repositories of the same type, which developers are most familiar with. In addition, local externalities have probabilistic, statistical, and computational advantages. First, local externalities ensure desirable properties of models for large networks (Schweinberger and Handcock, 2015). Second, local externalities facilitate theoretical guarantees for statistical procedures, which we discuss in Section 4. Third, local externalities enable local computing on subnetworks, which facilitates parallel computing and hence large-scale computing (Babkin et al., 2020).
@@ 3.2 Equilibrium
Software development is a dynamic process. We assume that time is discrete and that at each time point $t\in\{1,2,\ldots\}$ a single repository is updated. In other words, maintainers do not update software repositories continuously, but certain events – unobservable to researchers – trigger the need for an update. For example, if repository $i$ depends on repository $j$, then an update of repository $j$ may require an update of repository $i$, or a request by users of repository $i$ may trigger an update of repository $i$. The update can be produced in-house or by forming a dependency to some other repository $j$ containing the same functionality. Since researchers may not be able to observe the timing of such updates, we model updates as a discrete-time Markov process (Norris, 1997).
@@
The first ingredient of the discrete-time Markov process is the update process that determines which repository $i$ requires an update and which dependency $j$ is considered to implement the update. We assume that the probability that repository $i$ requires an update and the developers of repository $i$ consider creating, maintaining, or removing a dependence to repository $j$ can be expressed as a function $\rho_{i,j}$. A simple example is an update process that selects an ordered pair of repositories $i$ and $j$ at random. In other words, a repository $i$ is selected with probability $1/n$ and requires an update. Conditional on the event that repository $i$ requires an update, the developers of repository $i$ consider updating the possible dependency of repository $i$ on repository $j$ with probability $1/(n-1)$. As a result, the probability that repository $i$ requires an update and the developers of repository $i$ consider updating a (possible) dependency of repository $i$ on repository $j$ is
$\rho_{i,j}\left(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)\;\;=\;\;\dfrac{1}{n\,(n-1)},$
where $\bm{g}^{t-1}$ is the network at time $t-1$. A second example is an update process that allows the conditional probability that an update of the (possible) dependency of repository $i$ on repository $j$ is proposed to depend on the distance $d(\bm{x}_{i},\bm{x}_{j})$ between the repositories $i$ and $j$. The distance $d(\bm{x}_{i},\bm{x}_{j})$ is any distance function satisfying reflexity, symmetry, and the triangle inequality: e.g., $d(\bm{x}_{i},\bm{x}_{j})$ may be the Euclidean distance between observable characteristics $\bm{x}_{i}$ and $\bm{x}_{j}$ of repositories $i$ and $j$, that is, $d(\bm{x}_{i},\bm{x}_{j})\coloneqq|\!|\bm{x}_{i}-\bm{x}_{j}|\!|_{2}$. In other words, a repository $i$ is selected with probability $1/n$ and, conditional on the event that repository $i$ requires an update, the developers of repository $i$ consider updating the possible dependency of repository $i$ on repository $j$ with probability $e^{-d(\bm{x}_{i},\bm{x}_{j})}/\sum_{k=1}^{n}e^{-d(\bm{x}_{i},\bm{x}_{k})}$, which implies that
$\rho_{i,j}\left(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)\;\;=\;\;\dfrac{1}{n}\;\dfrac{e^{-d(\bm{x}_{i},\bm{x}_{j})}}{\sum_{k=1}^{n}e^{-d(\bm{x}_{i},\bm{x}_{k})}}.$
To characterize the stationary distribution of the discrete-time Markov process, the update process needs to satisfy two mild conditions (Mele, 2017). First, the update probabilities $\rho_{i,j}$ can depend on the entire software dependency network excluding $(i,j)$:
$\rho_{i,j}\Big{(}\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\Big{)}\;\;=\;\;\rho_{i,j}\Big{(}\bm{g}_{-(i,j)}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\Big{)}.$
Second, the update probabilities $\rho_{i,j}$ are strictly positive for all ordered pairs of repositories $(i,j)$:
@@
This implies that each possible dependency can be proposed with a positive probability, no matter how small. These assumptions help ensure that in the long-run equilibrium these probabilities do not affect the distribution of networks (Mele, 2017).
While the update process is stochastic, the decision to create, maintain, or remove a possible dependency of repository $i$ on repository $j$ is a strategic decision. Conditional on the opportunity to update the dependency of repository $i$ on repository $j$, the developers of repository $i$ make a decision about whether to create, maintain, or remove a dependency on repository $j$ based on utility function $U_{i}$. Before the developers of repository $i$ make a decision about the (possible) dependency on repository $j$, the developers receive random shocks $\varepsilon_{ij0}\in\mathbb{R}$ and $\varepsilon_{ij1}\in\mathbb{R}$ to the perceived utility of a dependency of repository $i$ on repository $j$ ($g_{ij}^{t}=1$) and the perceived utility of writing code in-house ($g_{ij}^{t}=0$). After receiving random shocks $\varepsilon_{ij0}\in\mathbb{R}$ and $\varepsilon_{ij1}\in\mathbb{R}$, the developers of repository $i$ create a dependency on repository $j$ if the perceived utility of the dependency of repository $i$ on repository $j$ ($g_{ij}^{t}=1$) exceeds the utility of writing code in-house ($g_{ij}^{t}=0$):
@@
where $\bm{g}_{-(i,j)}^{t-1}$ refers to network $\bm{g}^{t-1}$ excluding the possible dependency $g_{ij}$ of repository $i$ on repository $j$. The random shocks to payoffs $\varepsilon_{ij0}^{t}\in\mathbb{R}$ and $\varepsilon_{ij1}^{t}\in\mathbb{R}$ capture the fact that unexpected events can occur when developing code in-house ($\varepsilon_{ij0}$) and when creating a dependency ($\varepsilon_{ij1}$), e.g., changes in the composition of the developer team or additional information about repository $j$. We assume that the random shocks to payoffs $\varepsilon_{ij0}$ and $\varepsilon_{ij1}$ do not depend on time, are independent across pairs of repositories $(i,j)$, and follow a logistic distribution, which is a standard assumption in the literature on discrete choice models (Mele, 2017; Graham and dePaula, 2020; Boucher and Mourifie, 2017). As a consequence, when the opportunity to update the dependency of repository $i$ on repository $j$ arises, the conditional choice probability of a dependency of repository $i$ on repository $j$ is
$\mathbb{P}\left(g_{ij}^{t}=1\ |\ \bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta}\right)=\frac{e^{U_{i}(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})-U_{i}(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})}}{1+e^{U_{i}(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})-U_{i}(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})}}.$ (9)